# FOF-X: Towards Real-time Detailed Human Reconstruction from a Single Image

### v1

@@ IV-E4 Improved Inter-Conversion Algorithm
IV-F The Real-time Pipeline
To perform real-time monocular human reconstruction, we design a four-stage pipeline. In the first stage, we get an image from the video stream with OpenCV [55], use RVM [56] to get a human mask to remove the background, and predict a SMPL model as prior with 4D-humans[57]. RVM is originally a matting method, and we use a threshold of 0.5 to obtain a binary mask. In the second stage, a CUDA-implemented rasterizer is used to render dual-sided normal maps and a FOF representation for the predicted SMPL. The proposed mesh-to-FOF conversion algorithm is integrated into the rasterizer. In the third stage, the pre-trained dual-sided normal estimator and our reconstruction network are encapsulated into a single neural network module, which is used to reconstruct the FOF. In the final stage, the FOF is resized to a proper resolution ($256\times 256$ in our implementation) and transformed into a mesh. All four stages are implemented in PyTorch and accelerated with TensorRT, running on a single RTX 4090 GPU. Our pipeline works at 30+ FPS, which is limited by the camera (Logic C920 Pro). With a recorded video, it can achieve a frame rate of about 50 FPS. Fig. 11 shows the real-time reconstruction system. Please check the demo video for more results.
IV-G Application to the Reconstruction of General Objects
We apply our method to cars to further demonstrate the generalizability of our FOF representation. We use ShapeNet Core V2[58] and filter the dataset to remove data with corrupted geometry. However, the remaining meshes are still non-watertight and need further post-processing. We render the mesh into normal maps and depth maps to get an oriented point cloud from 128 views with Fibonacci lattice sampling on the sphere surface. Then we apply Poisson reconstruction to the point cloud to get regularized meshes. We only use RGB images to train the reconstruction model because of the lack of parametric models for cars. We collect 32 views of 1350 objects for training and 50 objects for testing. As illustrated in Fig. 12, although some artifacts can be seen on the wheels, the representation of FOF still produces promising results. This verifies that FOF has the potential to generalize to general objects and can be used for scene representations with proper modification.
Figure 12: FOF-X demonstrates strong generalization on other 3D objects. Here we show some reconstructed results on cars from the ShapeNet dataset.
@@ V Conclusion and discussion
V-A Conclusion
In this paper, we first present the Fourier Occupancy Field (FOF) representation for monocular real-time 3D human reconstruction and further extend it to the FOF-X framework for more detailed results. FOF is an efficient and flexible 3D representation defined as a 2D vector field orthogonal to the $z$-axis. At the core of FOF, we leverage the Fourier series as basis functions to perform subspace approximation of the occupancy field. Specifically, the occupancy field is expanded into the Fourier series along the $z$-axis; only the first few terms are used to describe the whole field. Each term corresponds to a single-channel 2D map. As a result, FOF can represent a 3D object in the form of a multi-channel image, which is pixel-aligned with an orthogonal image of the object. Thus, FOF can be estimated using a simple image-to-image network with an RGB image as input. FOF can be seamlessly inter-converted to and from the mesh representation, making it compatible with traditional pipelines. Additionally, as a compact and computationally efficient 3D representation, FOF is suitable for real-time applications like 3DTV.
When extending the FOF representation to the FOF-X reconstruction framework, we introduce additional improvements to the conversion algorithm and the reconstruction algorithm. For the conversion algorithm, an automaton-based discontinuity matcher is designed to convert a triangle mesh to a FOF efficiently and robustly. Thanks to this, the SMPL mesh prior can be added into the real-time pipeline, further enhancing the performance of the system. Besides, the FOF to mesh conversion is also improved with the Laplacian coordinate constraint, eliminating the stair-step artifacts. For the reconstruction framework, we predict dual-sided normal maps from input RGB images first and then use dual-sided normal maps instead of RGB images as the input of the main network, which greatly mitigates the performance degradation caused by the domain gap between the training images and the real images. The efficacy of FOF and FOF-X is validated on 3D scanned and in-the-wild datasets, where our approaches effectively recover detailed human geometry with better quality than previous solutions. Additional experiments on general objects showcase the generalizability of our method.
V-B Limitations and Future Work
The FOF representation is based on the Fourier series expansion of the square-wave-like function. When applying FOF to thin objects, the spectrum of the function will contain many high-frequency components. In such cases, we need more terms in the Fourier series to approximate the function. The number of terms we need is roughly inversely proportional to the thickness of the object. In our experiments, we found that FOF fails to reconstruct objects that are too thin, like detailed hands. Although wavelet transformation and short-time Fourier transformation seem to be a promising option, the implementation and computational efficiency still need more exploration. We leave this issue for future work to further improve our framework.
@@ VI Acknowledgements
This work was partially supported by the National Key R&D Program of China (2023YFC3082100) and the National Natural Science Foundation of China (62122058, 62125107, 62171317, and 62231018).

### v2

@@ V Conclusion and discussion
Conclusion. In this paper, we first present the Fourier Occupancy Field (FOF) representation for monocular real-time 3D human reconstruction and further extend it to the FOF-X framework for more detailed results. FOF is an efficient and flexible 3D representation defined as a 2D vector field orthogonal to the $z$-axis. When extending the FOF representation to the FOF-X reconstruction framework, we introduce additional improvements to the conversion algorithm and the reconstruction algorithm. For the conversion algorithm, an automaton-based discontinuity matcher is designed to convert a triangle mesh to a FOF efficiently and robustly. Thanks to this, the SMPL mesh prior can be added into the real-time pipeline, further enhancing the performance of the system. Besides, the FOF to mesh conversion is also improved with the Laplacian coordinate constraint, eliminating the stair-step artifacts. For the reconstruction framework, we predict dual-sided normal maps from input RGB images first and then use dual-sided normal maps instead of RGB images as the input of the main network, which greatly mitigates the performance degradation caused by the domain gap between the training images and the real images. The efficacy of FOF and FOF-X is validated on 3D scanned and in-the-wild datasets, where our approaches effectively recover detailed human geometry with better quality than previous solutions.
Limitations and Future Work. The FOF representation is based on the Fourier series expansion of the square-wave-like function. When applying FOF to thin objects, the spectrum of the function will contain many high-frequency components. In such cases, we need more terms in the Fourier series to approximate the function. The number of terms we need is roughly inversely proportional to the thickness of the object. In our experiments, we found that FOF fails to reconstruct objects that are too thin, like single layer mesh. Although wavelet transformation and short-time Fourier transformation seem to be a promising option, the implementation and computational efficiency still need more exploration. In addition, the FOF representation shows significant potential for broader applications. Its mathematical formulation and efficient parameterization could be extended to general object representation and even scene-level modeling through appropriate architectural modifications. These promising directions will be explored in future work to enhance the framework’s capabilities and applicability.
@@ VI Acknowledgements
This work was partially supported by the National Key R&D Program of China (2023YFC3082100) and the National Natural Science Foundation of China (62171317, 62125107, and 62231018).